OpenStudy (anonymous):

Use the Pythagorean theorem to find the hypotenuse if one leg is 34 and the other leg is 18.

OpenStudy (anonymous):

28.84 38.47 35.56 28.48 ?

OpenStudy (nubeer):

hyp^2=base^2+perpendicular^2 hyp^2=34^2+18^2 hyp^2=1480 hyp=38.47

OpenStudy (anonymous):

Thanks! And thank you for showing how to actually solve it, too.

OpenStudy (nubeer):

welcome :)